solve the given system of differential equations:

dx/dt=2x+2y

dy/dt=x+3


1
Expert's answer
2021-02-24T07:04:06-0500

solve the given system of differential equations:

"\begin{cases}
 \frac{dx}{dt}=2x+2y \\
 \frac{dy}{dt}=x+3 
\end{cases}"

Solution:

From the second equation:

"x=\frac{dy}{dt}-3"  (*)

"\frac{dx}{dt}=\frac{d^2y}{dt^2}"

Substitute "x" and "\frac{dx}{dt}" into the first equation:

"\frac{d^2y}{dt^2}=2(\frac{dy}{dt}-3 )+2y"

"\frac{d^2y}{dt^2}-2\frac{dy}{dt}-2y=-6" (**)

First find the general solution of the corresponding homogeneous equation:

"\frac{d^2y_0}{dt^2}-2\frac{dy_0}{dt}-2y_0=0"

Let's compose and solve the characteristic equation:

"\lambda^2-2\lambda-2=0"

"\lambda_1=1-\sqrt3" ,  "\lambda_2=1+\sqrt3" .

"y_0=C_1e^{\lambda_1t}+C_2e^{\lambda_2t}=C_1e^{(1-\sqrt3)t}+C_2e^{(1+\sqrt3)t}" .

Find a typical (specific) solution of the non-homogeneous equation in the form "y_p=A"

Derivatives of the solution: "\frac{dy_p}{dt}=0" , "\frac{d^2y_p}{dt^2}=0"

Substitute those "solutions" into the (**):

"-2A=-6"

"A=3"

Add the typical and the complementary solutions to get the complete solution:

"y=y_0+y_p=C_1e^{(1-\sqrt3)t}+C_2e^{(1+\sqrt3)t}+3" .

To find "x" substitute "\frac{dy}{dt}" into the (*):

"x=\frac{dy}{dt}-3 =C_1(1-\sqrt3)e^{(1-\sqrt3)t}+C_2(1+\sqrt3)e^{(1+\sqrt3)t}-3"

Answer: "\begin{cases}
 x =C_1(1-\sqrt3)e^{(1-\sqrt3)t}+C_2(1+\sqrt3)e^{(1+\sqrt3)t}-3 \\
 y=C_1e^{(1-\sqrt3)t}+C_2e^{(1+\sqrt3)t}+3
\end{cases}"